# Data Dependency Inference for Industrial Code Generation Based on UML Sequence Diagrams

### v2

@@ Abstract
Index Terms:
Data Dependency Inference, UML, sequence diagram, code generation
@@ I Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in automated code synthesis from natural language (NL) descriptions, transforming how developers approach programming tasks [1, 2, 3, 4, 5, 6]. However, when scaling from simple demonstrations to production-grade software systems, a fundamental challenge persists: the inherent imprecision of natural language becomes increasingly problematic for capturing the detailed specifications required by complex applications. Natural language descriptions often lack the structural clarity needed to unambiguously define sophisticated system interactions, execution flows, and data relationships[7, 8].
Contemporary software engineering practice increasingly adopts Unified Modeling Language (UML) sequence diagrams as design blueprints[9]. These diagrams provide formal visual syntax that explicitly captures control flows and service interactions with a obvious advancement over NL specifications, as demonstrated in model-driven engineering studies [10]. Recent work [11, 12] has systematically investigated LLM-based code generation from UML sequence diagrams, revealing persistent underperformance in industrial applications. The root cause lies in sequence diagrams’ inherent limitation: while effectively modeling control logic, they lack explicit representation of data dependencies, i.e., the lifeblood of operational execution. This forces LLMs to simultaneously reconstruct control flow and infer implicit data dependencies during code generation, a dual cognitive load that obviously amplifies the difficulty and error likelihood of code generation. Particularly for large-scale sequence diagrams or exceptionally intricate logic, LLMs are highly susceptible to generating severe hallucinations when inferring data dependencies.
@@ II-A Enhanced Sequence Diagram
Traditional methods for automatically generating code based on UML models [13, 14] exhibit significant limitations when faced with complex software systems. The main reason is that classic sequence diagrams are unable to fully express complex business processes and data dependencies, resulting in generated code that only covers simple scenarios and fails to meet the requirements of industrial complex systems. In response to the unique characteristics of complex software, we propose an enhanced sequence diagram design specification to better support subsequent data dependency inference and high-quality code generation.
@@ III-C Model and hyper-parameter
Due to confidentiality requirements regarding internal business data, we select the open-source models deepseek-r1-0528[15] and deepseek-v3-0324[16] for our experiments and deploy them on our own infrastructure. For all experiments, we set the temperature parameter to 0.1 to ensure more deterministic and stable outputs.
@@ VI-A UML Modeling and Analysis
The Unified Modeling Language (UML) is a standardized, general-purpose modeling language designed to represent the interactions among collaborating objects in software systems [17, 18, 19]. Over the past two decades, a variety of approaches leveraging UML for modeling and analysis tasks have been developed.
Early work focused on reverse engineering, which can be categorized into static and dynamic approaches[20, 21]. A seminal contribution in static analysis is the study by Rountev et al [22], which presents an algorithm for mapping reducible, exception-free control-flow graphs to UML interaction fragments. Conversely, dynamic approaches [23, 24, 25, 26] primarily focus on analyzing application performance through execution traces to generate sequence diagrams. These methods often utilize runtime data to reconstruct behavioral models that reflect actual system operations.
Reverse engineering reconstructs UML from existing code. Subsequent innovations have introduced more sophisticated frameworks. Notably, aToucan [27] proposes a rule-driven end-to-end system capable of automatically generating UML analysis models from use case descriptions written in constrained natural language. Similarly, Jahan et al. [28] propose an automated method to generate UML sequence diagrams from textual use cases.
Recent studies have further explored the application of LLMs. Ferrari et al. [29] systematically evaluate the capabilities of GPT-3.5 and GPT-4 in generating UML sequence diagrams directly from natural language requirements. This study highlights the potential of LLMs in UML modeling, demonstrating their effectiveness in translating textual descriptions into structured UML diagrams. However, it also identifies a critical gap in addressing Data Dependency Inference(DDI), a key issue that our paper addresses.
@@ VI-B From Design to Code: Challenges and Industrial Reality
Early model-driven approaches like UJECTOR [13] established foundational workflows for structural code generation from UML class diagrams and basic behavioral patterns from sequence/activity diagrams. Subsequent frameworks [14] improved standardization through XMI parsing and modular rule mapping. These rule-based systems typically require laborious manual adaptation when handling evolving design patterns.
The emergence of LLMs has introduced new paradigms. Sadik et al[12] leverages GPT-4 to bridge UML/OCL specifications with code generation, demonstrating enhanced adaptability compared to traditional template-based approaches. Recent multimodal approaches [30] attempt to address visual design artifacts through image-based UML processing.

### v3

# Data Dependency-Aware Code Generation from Enhanced UML Sequence Diagrams

@@ I Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in automated code synthesis from natural language (NL) descriptions, transforming how developers approach programming tasks [codex, hui2024qwen2, guo2024deepseek, JimenezYWYPPN24, codellama, gpt4o]. However, when scaling from simple demonstrations to production-grade software systems, a fundamental challenge persists: the inherent imprecision of natural language becomes increasingly problematic for capturing the detailed specifications required by complex applications. Natural language descriptions often lack the structural clarity needed to unambiguously define sophisticated system interactions, execution flows, and data relationships[DBLP:journals/tse/FerrariGST22, DBLP:conf/birthday/GervasiFZS19].
Contemporary software engineering practice increasingly adopts Unified Modeling Language (UML) sequence diagrams as design blueprints[shaikh2021more]. These diagrams provide formal visual syntax that explicitly captures control flows and service interactions with a obvious advancement over NL specifications, as demonstrated in model-driven engineering studies [DBLP:conf/models/MussbacherABBKMWW14]. Recent work [DBLP:conf/llm4code/AntalVF24, DBLP:conf/modelsward/SadikBO24] has systematically investigated LLM-based code generation from UML sequence diagrams, revealing persistent underperformance in industrial applications. The root cause lies in sequence diagrams’ inherent limitation: while effectively modeling control logic, they lack explicit representation of data dependencies, i.e., the lifeblood of operational execution. This forces LLMs to simultaneously reconstruct control flow and infer implicit data dependencies during code generation, a dual cognitive load that obviously amplifies the difficulty and error likelihood of code generation. Particularly for large-scale sequence diagrams or exceptionally intricate logic, LLMs are highly susceptible to generating severe hallucinations when inferring data dependencies.
@@ II-A Enhanced Sequence Diagram
Traditional methods for automatically generating code based on UML models [usman2008ujector, kluisritrakul2016generation] exhibit significant limitations when faced with complex software systems. The main reason is that classic sequence diagrams are unable to fully express complex business processes and data dependencies, resulting in generated code that only covers simple scenarios and fails to meet the requirements of industrial complex systems. In response to the unique characteristics of complex software, we propose an enhanced sequence diagram design specification to better support subsequent data dependency inference and high-quality code generation.
@@ III-C Model and hyper-parameter
Due to confidentiality requirements regarding internal business data, we select the open-source models deepseek-r1-0528[deepseekr1] and deepseek-v3-0324[deepseekv3] for our experiments and deploy them on our own infrastructure. For all experiments, we set the temperature parameter to 0.1 to ensure more deterministic and stable outputs.
@@ VI-A UML Modeling and Analysis
The Unified Modeling Language (UML) is a standardized, general-purpose modeling language designed to represent the interactions among collaborating objects in software systems [booch1996unified, ozkaya2020survey, hammond2006tahuti]. Over the past two decades, a variety of approaches leveraging UML for modeling and analysis tasks have been developed.
Early work focused on reverse engineering, which can be categorized into static and dynamic approaches[briand2006toward, abidi2015new]. A seminal contribution in static analysis is the study by Rountev et al [rountev2005static], which presents an algorithm for mapping reducible, exception-free control-flow graphs to UML interaction fragments. Conversely, dynamic approaches [delamare2006reverse, sarkar2013reverse, richner2002using, oechsle2002javavis] primarily focus on analyzing application performance through execution traces to generate sequence diagrams. These methods often utilize runtime data to reconstruct behavioral models that reflect actual system operations.
Reverse engineering reconstructs UML from existing code. Subsequent innovations have introduced more sophisticated frameworks. Notably, aToucan [aToucan] proposes a rule-driven end-to-end system capable of automatically generating UML analysis models from use case descriptions written in constrained natural language. Similarly, Jahan et al. [jahan2021generating] propose an automated method to generate UML sequence diagrams from textual use cases.
Recent studies have further explored the application of LLMs. Ferrari et al. [FerrariAA24] systematically evaluate the capabilities of GPT-3.5 and GPT-4 in generating UML sequence diagrams directly from natural language requirements. However, it also identifies a critical gap in addressing Data Dependency Inference(DDI), a key issue that our paper addresses.
@@ VI-B From Design to Code: Challenges and Industrial Reality
Early model-driven approaches like UJECTOR [usman2008ujector] established foundational workflows for structural code generation from UML class diagrams and basic behavioral patterns from sequence/activity diagrams. Subsequent frameworks [kluisritrakul2016generation] improved standardization through XMI parsing and modular rule mapping.
The emergence of LLMs has introduced new paradigms. Sadik et al[DBLP:conf/modelsward/SadikBO24] leverages GPT-4 to bridge UML/OCL specifications with code generation, demonstrating enhanced adaptability compared to traditional template-based approaches. Recent multimodal approaches [bates2025unified] attempt to address visual design artifacts through image-based UML processing.
@@ VII CONCLUSION
Acknowledgments
We appreciate the assistance from the WeChat pay team for their valuable contributions. This research is supported by the National Natural Science Foundation of China under project (No. 62472126, 62276075, 62192731), Natural Science Foundation of Guangdong Province (Project No. 2023A1515011959), Shenzhen-Hong Kong Jointly Funded Project (Category A, No. SGDX20230116 091246007) and the Tencent WeChat Rhino-Bird Focused Research Program.